OpenStudy (anonymous):

Use either substitution or elimination to solve 3x-7y=5 x+3y=-1 I used elimination..and got it wrong..I don't see how you can use two different ways of solving and get the same answer...Please explain..Thanks!

jimthompson5910 (jim_thompson5910):

x+3y=-1 x=-1-3y -------------------- 3x-7y=5 3(-1-3y)-7y=5 -3-9y-7y=5 -16y-3=5 -16y=5+3 -16y=8 y=8/(-16) y=-1/2 ---------------------- x=-1-3y x=-1-3(-1/2) x=-1+3/2 x=-2/2+3/2 x=(-2+3)/2 x=1/2 So the solutions are x=1/2 and y=-1/2

OpenStudy (anonymous):

Thanks! but that's substitution, what about elimination?

jimthompson5910 (jim_thompson5910):

3x-7y=5 x+3y=-1 3x-7y=5 -3(x+3y)=-3(-1) 3x-7y=5 -3x-9y=3 -16y=8 y=8/(-16) y=-1/2 ------------------ x+3y=-1 x+3(-1/2)=-1 x-3/2=-1 x=-1+3/2 x=1/2

OpenStudy (anonymous):

Thanks I understand it so much better now!!(:

jimthompson5910 (jim_thompson5910):

yw